# Paired Completion: Flexible Quantification of Issue-framing at Scale with LLMs

### v1

@@ Abstract
Detecting and quantifying issue framing in textual discourse - the perspective one takes to a given topic (e.g. climate science vs. denialism, misogyny vs. gender equality) - is highly valuable to a range of end-users from social and political scientists to program evaluators and policy analysts. However, conceptual framing is notoriously challenging for automated natural language processing (NLP) methods since the words and phrases used by either ‘side’ of an issue are often held in common, with only subtle stylistic flourishes separating their use. Here we develop and rigorously evaluate new detection methods for issue framing and narrative analysis within large text datasets. By introducing a novel application of next-token log probabilities derived from generative large language models (LLMs) we show that issue framing can be reliably and efficiently detected in large corpora with only a few examples of either perspective on a given issue, a method we call ‘paired completion’. Through 192 independent experiments over three novel, synthetic datasets, we evaluate paired completion against prompt-based LLM methods and labelled methods using traditional NLP and recent LLM contextual embeddings. We additionally conduct a cost-based analysis to mark out the feasible set of performant methods at production-level scales, and a model bias analysis. Together, our work demonstrates a feasible path to scalable, accurate and low-bias issue-framing in large corpora.
@@ 1 Introduction
It is widely held that public narratives have the power – for better and worse – to shape society (Shiller 2019; Patterson and Monroe 1998; Graber 2002; Barabas and Jerit 2009). For quantitative social scientists, a typical analytical strategy to quantify the occurrence, characteristics and dynamics of these important narratives is to use a ‘framing’ lens. According to the much-cited definition found in Entman (1993), framing is the process by which individuals “select some aspects of a perceived reality and make them more salient in a communicating text” with the purpose of promoting a particular interpretation or evaluation of reality. In essence, to frame, is to impose a world-view or ‘way-of-thinking’ in communication, with the hope that others will be persuaded to be convinced of the same. In the standard approach to framing analysis (Chong and Druckman 2007), one begins first, by identifying an issue (e.g. ‘climate change’); second, by defining the dimensions of that issue (e.g. ‘causes’, ‘economic impact’, etc.); third, by developing framings of those dimensions (e.g. climate change/ causes/ framing: ‘anthropogenic emissions are responsible for climate change’); and then finally, by the manual labelling of texts (sentences, paragraphs) as to their framing alignment. Computational approaches to framing quantification have addressed various aspects of this decomposition, often (unhelpfully) under the generic heading of ‘framing’ (we return to this point below) (Ali and Hassan 2022).
Almost all prior approaches to computational framing analysis consider the task as a supervised machine-learning problem, typically as a multi-class classification task (Field et al. 2018), and most often focusing on automatic labelling of dimensions (e.g. ‘2nd Amendment’, ‘Politics’, ‘Public Opinion’) of a single issue (e.g. gun violence) (Liu et al. 2019a; Zhang et al. 2023), rather the more elaborate ‘world-view’ like conceptualisation that Entman (1993) and Chong and Druckman (2007) hold. Where studies consider conceptual framing identification, large amounts of labelled data are required, and reported accuracy is very low (below 0.6) (Morstatter et al. 2018; Mendelsohn, Budak, and Jurgens 2021), demonstrating the severe challenges inherent in automating an already difficult human-level task.
@@
Paired completion takes advantage of the log-probability (logprob) outputs of an LLM [^1] Note: logprobs are available as outputs through the OpenAI API for “babbage-002” and “davinci-002” (OpenAI 2024a), and can be gathered by running the “vLLM OpenAI-compatible API” (Kwon et al. 2023) on a local machine, for a wide variety of open-source models. to find conditional probabilities of a text given a series of conditioners from different conditioning sets. We use the relative differences in probabilities to establish a baseline metric that (at least theoretically) is resilient to the model’s prior probabilities of both the conditioning text and the text being aligned to the conditioning sets. We demonstrate empirically that this method is successful, and that one achieves superior performance from using this method with raw base models compared to “asking” instruct-fine-tuned AI-models the “question” at hand.
We conduct rigorous evaluation of our proposed method across 192 independent experiments by comparing it to four framing classification approaches over four diverse, synthetic textual datasets, including two baseline approaches (traditional tf-idf vectors (Sparck Jones 1972; Salton 1983), and fasttext sentence embeddings (Bojanowski et al. 2017)) and three LLM-based methods (contextual embeddings (Peters et al. 2018; Devlin et al. 2018; OpenAI 2024b) LLM chat token probabilities (Radford et al. 2019), and our novel paired completion method). We demonstrate that the LLM-based approaches are, in general, far superior to the alternatives. The LLM-embedding approach is powerful with enough training data, but with small amounts of data (e.g. five sentences for each conditioning set) the LLM methods easily outperform LLM-embeddings. We also demonstrate that paired completion with LLMs is generally superior to the LLM prompting approach. We discuss why this might be the case in Section 2.1, from a theoretical perspective. We also conduct cost- and bias- comparison analysese at current gated API pricing to assess any trade-offs in performance.
@@ 1.1 Contributions
We introduce paired completion as a solution to the problem of textual alignment. We construct a series of high-quality synthetic datasets using a novel method which captures nuances of discourse on complex topics, and use these datasets to evaluate the performance of several approaches to textual alignment. We demonstrate that paired completion is a novel, efficient, and more effective method for performing textual alignment (compared to a chat-based LLM baseline).
@@ 1.2 Related Literature
Unfortunately, ‘framing’ analysis does not have a clear definition in the computational literature, as evidenced by the variety of tasks that arise in a recent survey of 37 ‘framing’ studies (Ali and Hassan 2022). A starting point for many framing approaches is to leverage existing corpora of labelled datasets. Here, the media frames corpus (MFC) (Card et al. 2015) and the gun violence frame corpus (GVFC) (Liu et al. 2019a) have been the basis of many methodological contributions. However, these datasets conceptualise ‘framing’ as dimensions (ala Chong and Druckman (2007)) of a topic or issue, not conceptual frames as we have distinguished earlier. The MFC is composed of 15 generic ‘frames’ (Boydstun et al. 2013) such as ‘economic’, ‘public opinion’ and ‘cultural identity’ applied to three issues (‘immigration’, ‘smoking’, and ‘same-sex marriage’). Thousands of annotations were recorded as to whether one of these dimensions were associated with the text on a given issue. Likewise, the GVFC follows a similar approach, albeit tied more tightly to gun violence, ‘frames’ are equivalent to dimensions, and include ‘2nd Amendment’, ‘Politics’, and ‘Public Opinion’. Typically, computational methods approach framing in this way as a multi-class classification problem, using supervised machine learning methodologies such as featuring engineering, classifier selection and k-fold evaluation (Field et al. 2018; Liu et al. 2019a; Akyürek et al. 2020; Zhang et al. 2023). Common to all of these approaches is the need for large amounts of labelled ground-truth data, and consequently, the outcome that methods are not generalisable beyond the topics under study.
@@
Where ‘framing’ is implemented in a closer way to the conceptual framing we address in this work, challenges remain around the need for large labelled datasets and the accuracy of the methods. Morstatter et al. (2018) consider support for, or against, 10 framings related to Balistic Missile Defence (BMD) in Europe over 823 online news articles (31k sentences). By writing out support- and opposition- (polarity) perspectives for each of the 10 framings, they are able to generate 20 framing-polarity classes, and apply traditional NLP methods to multi-class prediction. Alternatively, Mendelsohn et al. (2021) label 3.6k social media posts (tweets) related to immigration for the 15 generic ‘framings’ of Boydstun et al. (Boydstun et al. 2013) together with 11 conceptual framings (e.g. ‘hero’, ‘victim’, ‘threat’ positions on immigrants). Using a base and fine-tuned encoder-only transformer approach (Devlin et al. 2018; Liu et al. 2019b), they again conduct a multi-label classification study. Whilst these examples are closely aligned to the same conceptual framing identification problem we address in the current study, each requires thousands of hand-labelled data to develop features to train traditional supervised machine-learning algorithms. Furthermore, and underlining the challenge of this task for traditional (even transformer based encoder methods), accuracy scores across 20 polarity classes in Morstatter et al. and f1 scores across 11 conceptual framings in Mendelsohn, Budak, and Jurgens are up to just 0.434 and 0.552 respectively.
Alternatively, and closer to the spirit of the present work, Guo, Ma, and Vosoughi (2022) take on a related but distinct task of quantifying the similarity between news sources by fine-tuning LLMs to each source and then conducting differential experiments on the likelihoods of word-completions when masking specific words within sentences on specific topics. The approach is highly interesting and would enable the characterisation of how a source (e.g. a news outlet, a speaker, a social media account) conceptually framed a specific policy area relative to another source. However, fine-tuning a LLM requires a large amount of text from each source to be feasible, more than would be available for more than just the most prominent political- or business- leaders. And, whilst the method could be applied to sub-sets of the corpora by time-interval (reminiscent of computational cultural analysis (Kozlowski, Taddy, and Evans 2019)) this further reduces the texts available.
Together, existing computational approaches to ‘framing’, whether focusing on the different, higher-level task of dimension identification, or studying conceptual framing as we do in the present work, require large amounts of labelled data to train multi-class classifiers. In contrast, paired completion, studies in the present work, proceeds with minimal examples of the conceptual framings required (e.g. 5-10) and requires no ‘training’ as such.
Stance detection
@@
Finally, a related task is that of stance detection, which we address briefly. Stance detection is typically formalised as the, “automatic classification of the stance of the producer of a piece of text, towards a target, into one of three classes: Favor, Against, Neither.” (Küçük and Fazli 2020) In effect, stance detection is a sub-problem of sentiment analysis, and again, typical approaches leverage traditional NLP techniques with labelled data as inputs (Küçük and Fazli 2020). Whilst paired completion shares the notion of ‘target texts’, these texts work in concert to mark out a complex, nuanced conceptual framing on any issue, and the two priming sets (A,B) need not be strictly in opposition, but represent two perspectives, opening up more complex analytical insights than simply ‘favor’ or ‘against’.
Perplexity
@@
One common measure of the capability of an LLM is perplexity (Jelinek et al. 1977), which is a statistical measure of the model’s “surprise” at a given completion under the logic that a model which is less surprised by correct answers is better (similar to the maximum likelihood principle). The paired completion approach developed in this work is a measure similar to perplexity, but instead of seeking the estimated likelihood of a particular completion we instead calculate and compare the likelihoods of multiple completions of a given text.
@@ 2 Textual Alignment & Paired Completion
We conjecture that LLMs are well suited to performing the issue-framing task since, with the advent of attentional transformer technology, they have been shown to be remarkably successful at modelling human language. That is, forming a deep abstract representation of the meaning of human communication. Precisely the capability we require to assess textual alignment.
@@ 2.1 Paired Completion
For example, a prompt sequence could be ‘[priming text from A, $s_{1}$] Owning a dog will improve your life. [target text, $x$] Dog owners have lower blood pressure and less stress in general.’ A similar sequence would be created for the same text $x$ with priming text(s) from set B. Each prompt sequence is then passed, one at a time, to a generative LLM, and instead of seeking a completion (i.e. generating new tokens) from the LLM, we instead exploit many LLM’s ability to provide log-probabilities (the log of the likelihood that the model would have chosen that token/word next) for each token passed to the language model as if it had generated this exact sequence of text. By so doing, we generate two conditional log-probabilities, $lp(x|s_{1})$ and $lp(x|s_{2})$ (see details in sec 2.1), the conditional log-probs of $x$ being the completion to the priming sequence $s_{1}$ and $s_{2}$ respectively.
@@
To summarise, paired-completion leverages the ‘deep’ language modelling properties of LLM – their deep contextual representation of human meaning – to quantify the likelihood that a target text will follow from a given conditioning prior. As such, we conjecture that powerful LLMs that are neither fine-tuned (to a particular task or corpus) nor moderated by post-training methods to suppress some behaviours and up-promote others (e.g. reinforcement learning with human feedback, RHLF (Ziegler et al. 2020)) will be most well suited to the paired completion method. Either of these adaptations of LLMs could reasonably trade off fundamental (general) language modelling capabilities of LLMs in service of performance or safety on a particular task or in a particular context.
See the Supplementary Information for details of the implementation of this method in evaluation.
The Diff Metric
We define the diff metric as follows:
@@
Note that $\Delta(s_{1},s_{2},x)=-\Delta(s_{2},s_{1},x)$.
Since the core idea of paired completion is to use the priming/conditioning sequence to statistically deflect the LLM towards the given framing (and so, measure the model’s degree of ‘surprise’ with the completion text) we conjecture that a longer conditioning sequence may lead to improved accuracy in classifying and retrieving texts aligned with a given framing. To explore this possibility we test two treatments, with either one ($k=1$) or two ($k=2$) priming/conditioning text(s) being used. Implementation details are provided in Supplementary Information.
3 Synthetic Dataset Formation
As described earlier, existing ‘framing’ datasets, such as the MFC (Card et al. 2015) and GVFC (Liu et al. 2019a) are not well suited for application to the task we study here since they label dimensions of a topic as ‘frames’. We are not aware of another comparable dataset that makes labels of conceptual framing available across a number of issues. For this reason, conducting evaluation with synthetic data was considered for the present study, although acknowledging inherent limitations (see last section). That said, there are some positive attributes of using synthetic data that we briefly outline.
First, our initial experience with practitioners in fields that are attempting to change public narratives, demonstrated that non-synthetic (human authored) examples of framings can carry correlated linguistic features that may pollute analysis. For example, found narratives that carry a misogynistic perspective can be relative short and abrupt, whilst opposing narratives which speak for gender equality often are expressed with longer, more complex reasoning. Early testing showed that LLMs could pick up on linguistic features such as length and complexity, confusing the signal. Whereas, our paired synthetic pipeline (see full prompts etc. in Supplementary Information) is designed to provide a very balanced (tone, length, complexity etc.) dataset, with only the conceptual framing as the distinguishing feature of the texts.

### v2

@@ Abstract
Detecting issue framing in text - how different perspectives approach the same topic - is valuable for social science and policy analysis, yet challenging for automated methods due to subtle linguistic differences. We introduce ‘paired completion’, a novel approach using LLM next-token log probabilities to detect contrasting frames using minimal examples. Through extensive evaluation across synthetic datasets and a human-labeled corpus, we demonstrate that paired completion is a cost-efficient, low-bias alternative to both prompt-based and embedding-based methods, offering a scalable solution for analyzing issue framing in large text collections, especially suited to low-resource settings.
Paired Completion: Flexible Quantification of Issue-framing at Scale with LLMs
Simon D. Angus Dept. of Economics & SoDa Laboratories Monash Business School Monash University simon.angus@monash.edu                        Lachlan O’Neill SoDa Laboratories Monash Business School Monash University lachlan.oneill@monash.edu
@@ 1 Introduction
It is widely held that public narratives have the power – for better and worse – to shape society Shiller (2019); Patterson and Monroe (1998); Graber (2002); Barabas and Jerit (2009). For quantitative social scientists, a typical analytical strategy to quantify the occurrence, characteristics and dynamics of these important narratives is to use a ‘framing’ lens. According to the much-cited definition found in Entman Entman (1993), framing is the process by which individuals “select some aspects of a perceived reality and make them more salient in a communicating text” with the purpose of promoting a particular interpretation or evaluation of reality. In essence, to frame, is to impose a world-view or ‘way-of-thinking’ in communication, with the hope that others will be persuaded to be convinced of the same. In the standard approach to framing analysis Chong and Druckman (2007), one begins first, by identifying an issue (e.g. ‘climate change’); second, by defining the dimensions of that issue (e.g. ‘causes’, ‘economic impact’, etc.); third, by developing framings of those dimensions (e.g. climate change/ causes/ framing: ‘anthropogenic emissions are responsible for climate change’); and then finally, by the manual labelling of texts (sentences, paragraphs) as to their framing alignment. Computational approaches to framing quantification have addressed various aspects of this decomposition, often (unhelpfully) under the generic heading of ‘framing’ (we return to this point below) Ali and Hassan (2022).
Almost all prior approaches to computational framing analysis consider the task as a supervised machine-learning problem, typically as a multi-class classification task Field et al. (2018), and most often focusing on automatic labelling of dimensions (e.g. ‘2nd Amendment’, ‘Politics’, ‘Public Opinion’) of a single issue (e.g. gun violence) Liu et al. (2019a); Zhang et al. (2023), rather the more elaborate ‘world-view’ like conceptualisation that Entman Entman (1993) and Chong and Druckman Chong and Druckman (2007) hold. Where studies consider conceptual framing identification, large amounts of labelled data are required, and reported accuracy is modest (below 0.6) Morstatter et al. (2018); Mendelsohn et al. (2021a), demonstrating the severe challenges inherent in automating an already difficult human-level task.
@@
Paired completion takes advantage of the log-probability (logprob) outputs of an LLM [^1] Note: logprobs are available as outputs through the OpenAI API for “babbage-002” and “davinci-002” (OpenAI, 2024a), and can be gathered by running the “vLLM OpenAI-compatible API” (Kwon et al., 2023) on a local machine, for a wide variety of open-source models. to find conditional probabilities of a text given a series of conditioners from different conditioning sets. We use the relative differences in probabilities to establish a baseline metric that (at least theoretically) is resilient to the model’s prior probabilities of both the conditioning text and the text being aligned to the conditioning sets. We demonstrate empirically that this method is successful, and that one achieves superior performance from using this method with raw base models compared to “asking” instruct-fine-tuned AI-models the “question” at hand.
We conduct rigorous evaluation of our proposed method across a large synthetic dataset and human-labelled Immigration Tweet datasets. Our first study on synthetic data includes 192 independent experiments which compare paired-completion to four framing classification approaches over four diverse textual datasets, including two baseline approaches (traditional tf-idf vectors (Sparck Jones, 1972; Salton, 1983), and fasttext sentence embeddings (Bojanowski et al., 2017)) and three LLM-based methods (contextual embeddings (Peters et al., 2018; Devlin et al., 2018; OpenAI, 2024b) LLM chat token probabilities (Radford et al., 2019), and our novel paired completion method). We demonstrate that the LLM-based approaches are, in general, far superior to the alternatives. The LLM-embedding approach is powerful with enough training data, but with small amounts of data (e.g. five sentences for each conditioning set) the LLM methods easily outperform LLM-embeddings. We also demonstrate that paired completion with LLMs is generally superior to the LLM prompting approach. We discuss why this might be the case in Section B, from a theoretical perspective. We also conduct cost- and bias- comparison analysese at current gated API pricing to assess any trade-offs in performance. Our Immigration Tweet evaluation further demonstrates that paired-completion is strongly performant in diverse settings, especially when framing pairs under study are semantically contrasting.
@@ 1.1 Contributions
We introduce paired completion as a promising tool in the textual alignment task, especially in low-resource settings. We construct and make available a series of high-quality synthetic datasets. On both synthetic and real datasets we demonstrate that paired completion is a novel, efficient, low-bias alternative to either a chat-based LLM baseline or an embedding-training approach.
@@ 1.2 Related Literature
Unfortunately, ‘framing’ analysis does not have a clear definition in the computational literature, as evidenced by the variety of tasks that arise in a recent survey of 37 ‘framing’ studies Ali and Hassan (2022). A starting point for many framing approaches is to leverage existing corpora of labelled datasets. Here, the media frames corpus (MFC) Card et al. (2015) and the gun violence frame corpus (GVFC) Liu et al. (2019a) have been the basis of many methodological contributions. However, these datasets conceptualise ‘framing’ as dimensions (ala Chong and Druckman Chong and Druckman (2007)) of a topic or issue, not conceptual frames as we have distinguished earlier. The MFC is composed of 15 generic ‘frames’ Boydstun et al. (2013) such as ‘economic’, ‘public opinion’ and ‘cultural identity’ applied to three issues (‘immigration’, ‘smoking’, and ‘same-sex marriage’). Thousands of annotations were recorded as to whether one of these dimensions were associated with the text on a given issue. Likewise, the GVFC follows a similar approach, albeit tied more tightly to gun violence, ‘frames’ are equivalent to dimensions, and include ‘2nd Amendment’, ‘Politics’, and ‘Public Opinion’. Typically, computational methods approach framing in this way as a multi-class classification problem, using supervised machine learning methodologies such as featuring engineering, classifier selection and k-fold evaluation Field et al. (2018); Liu et al. (2019a); Akyürek et al. (2020); Zhang et al. (2023). Common to all of these approaches is the need for large amounts of labelled ground-truth data, and consequently, the outcome that methods are not generalisable beyond the topics under study.
@@
Where ‘framing’ is implemented in a closer way to the conceptual framing we address in this work, challenges remain around the need for large labelled datasets and the accuracy of the methods. Morstatter et al. Morstatter et al. (2018) consider support for, or against, 10 framings related to Balistic Missile Defence (BMD) in Europe over 823 online news articles (31k sentences). By writing out support- and opposition- (polarity) perspectives for each of the 10 framings, they are able to generate 20 framing-polarity classes, and apply traditional NLP methods to multi-class prediction. Alternatively, Mendelsohn et al. Mendelsohn et al. (2021a) label 3.6k social media posts (tweets) related to immigration for the 15 generic ‘framings’ of Boydstun et al. Boydstun et al. (2013) together with 11 conceptual framings (e.g. ‘hero’, ‘victim’, ‘threat’ positions on immigrants). Using a base and fine-tuned encoder-only transformer approach Devlin et al. (2018); Liu et al. (2019b), they again conduct a multi-label classification study (we return to this dataset below). Whilst these examples are closely aligned to the same conceptual framing identification problem we address in the current study, each requires thousands of hand-labelled data to develop features to train traditional supervised machine-learning algorithms. Furthermore, and underlining the challenge of this task for traditional (even transformer based encoder methods), accuracy scores across 20 polarity classes in Morstatter et al. and f1 scores across 11 conceptual framings in Mendelsohn et al. are up to just 0.434 and 0.552 respectively.
Alternatively, and closer to the spirit of the present work, Guo et al. Guo et al. (2022) take on a related but distinct task of quantifying the similarity between news sources by fine-tuning LLMs to each source and then conducting differential experiments on the likelihoods of word-completions when masking specific words within sentences on specific topics.
1.2.2 Frame discovery
A closely related task to textual alignment task is framing discovery. Here, working in an unsupervised or semi-supervised manner, the task is to identify prominent framings in a corpus of texts. Examples include Demszky et al. (2019) who cluster tweet embeddings, Roy and Goldwasser (2020) and Roy et al. (2021) who develop lexicons, and train embeddings to develop a frame identification model. The approach introduced in the present work complements these more elaborate discovery methods, especially in low-resource settings where advocacy groups are already aware of the ‘unhelpful’ and ‘helpful’ frames related to their domain, and so can move directly to known-framing detection.
1.2.3 Stance detection
@@
A related task is that of stance detection, which we address briefly. Stance detection is typically formalised as the, “automatic classification of the stance of the producer of a piece of text, towards a target, into one of three classes: Favor, Against, Neither.” Küçük and Fazli (2020) In effect, stance detection is a sub-problem of sentiment analysis, and again, typical approaches leverage traditional NLP techniques with labelled data as inputs Küçük and Fazli (2020). Whilst paired completion shares the notion of ‘target texts’, these texts work in concert to mark out a complex, nuanced conceptual framing on any issue, and the two priming sets (A,B) need not be strictly in opposition, but represent two perspectives, opening up more complex analytical insights than simply ‘favor’ or ‘against’.
1.2.4 Perplexity
@@
One common measure of the capability of an LLM is perplexity (Jelinek et al., 1977), which is a statistical measure of the model’s “surprise” at a given completion under the logic that a model which is less surprised by correct answers is better (similar to the maximum likelihood principle). The paired completion approach developed in this work is a measure similar to perplexity, but instead of seeking the estimated likelihood of a particular completion we instead calculate and compare the likelihoods of multiple completions of a given text.
@@ 2.1 Paired Completion
For example, a prompt sequence could be ‘[priming text from A, $s_{1}$] Owning a dog will improve your life. [target text, $x$] Dog owners have lower blood pressure and less stress in general.’ A similar sequence would be created for the same text $x$ with priming text(s) from set B. Each prompt sequence is then passed, one at a time, to a generative LLM, and instead of seeking a completion (i.e. generating new tokens) from the LLM, we instead exploit many LLM’s ability to provide log-probabilities (the log of the likelihood that the model would have chosen that token/word next) for each token passed to the language model as if it had generated this exact sequence of text. By so doing, we generate two conditional log-probabilities, $lp(x|s_{1})$ and $lp(x|s_{2})$ (see details in sec B), the conditional log-probs of $x$ being the completion to the priming sequence $s_{1}$ and $s_{2}$ respectively.
@@
Paired-completion leverages LLMs’ deep contextual representation of human meaning to evaluate text likelihood given prior context. We hypothesize that base LLMs, without task-specific fine-tuning or RLHF moderation, are best suited for this method, as such adaptations may compromise their fundamental language modeling capabilities.
See the appendices for details of the implementation of this method in evaluation.
2.2 The Diff Metric
To quantify textual alignment, we introduce the Diff metric which measures the relative likelihood of a text following from different conditioning priors. Given a target text $x$ and two conditioning texts $s_{1}$ and $s_{2}$, we compute:
@@ 2.2 The Diff Metric
where $\text{lp}(x|s)$ represents the conditional log-probability of text $x$ given conditioning text $s$. The metric captures whether text $x$ is more likely to follow $s_{1}$ or $s_{2}$ by comparing their respective conditional probabilities. A positive value indicates stronger alignment with $s_{1}$, while a negative value indicates stronger alignment with $s_{2}$. This approach is robust to the model’s prior probabilities of both the conditioning texts and the target text, as it considers only the relative difference in conditional probabilities. Detailed mathematical derivations and implementation specifics can be found in the appendices.